OpenStudy (anonymous):

find the sum of the first 90 terms of the sequence -4,-1,2,5,8... a. 11654 b.11657 c.11656 d.11655

OpenStudy (anonymous):

Since it is moving 3 by 3 we can assume \[u _{n}=3n + c _{1}\]\[u _{1}=-4 \rightarrow c _{1}=-7\] Becoming,\[u _{n}=3n-7\] \[\sum_{n=1}^{90}3n-7=(-7)\times90+3\sum_{n=1}^{90}n=-630+3\times90\times(90+1)\div2=11655\] Some other time I might explain why \[\sum_{k=1}^{n}k=n \times(n+1)\div2\] Anyway answer is d. :) Cheers PS: FIRST POST, BANZAI!!!!